# Reasoning Beyond Limits: Advances and Open Problems for LLMs

### v1

@@ III-C Mistral AI Small 3 24B
Mistral Small 3 [41] is a highly efficient language model, competing with Llama 3.3 70B and Qwen 32B, while being 3× faster on the same hardware. Optimized for 80% of generative AI tasks, it excels in instruction-following with low latency and achieves 81% accuracy on MMLU at 150 tokens/s. Fewer layers significantly reduce inference time, making it ideal for local deployment. Unlike models such as DeepSeek R1, Mistral Small 3 has not been trained using Reinforcement Learning (RL) or synthetic data, making it an earlier-stage model in the production pipeline with the potential for additional refinement.
@@ III-D DeepSeek-R1
The DeepSeek-R1 [42] series proposes advanced reasoning-optimized language models, including DeepSeek-R1-Zero, which is trained exclusively via Reinforcement Learning (RL) without Supervised Fine-Tuning (SFT), and DeepSeek-R1, which incorporates multi-stage training with cold-start data for improved reasoning and language coherence. DeepSeek-R1 matches OpenAI’s o1-1217 in reasoning tasks, leveraging key architectural innovations: Mixture of Experts (MoE) for efficient parameter utilization, Multihead Latent Attention (MLA) for reduced memory and compute costs, Multi-Token Prediction (MTP) for 2-3x faster token generation, and FP8 Quantization, which cuts memory usage by 75% while maintaining stability. DeepSeek has released DeepSeek-R1-Zero, DeepSeek-R1, and six dense models (1.5B–70B) to support open research, providing a scalable, high-performance foundation for AI-driven reasoning and inference.
@@
Qwen2.5-VL [43] is a significant upgrade to Qwen2-VL [63], enhancing vision-language capabilities across three sizes (3B, 7B, 72B). The Qwen2-VL Series proposes Naive Dynamic Resolution (NDR) for adaptive visual tokenization, enhancing efficiency and accuracy in vision-language tasks. It integrates Multimodal Rotary Position Embedding (M-RoPE) for improved text-image-video fusion and follows a unified image-video processing paradigm. By scaling both model size (2B, 8B, 72B) and training data, Qwen2-VL achieves competitive performance, with Qwen2-VL-72B matching GPT-4o and Claude 3.5-Sonnet on multimodal benchmarks. These innovations position Qwen2-VL as a state-of-the-art LVLM, advancing vision-language integration.
@@ III-F HuatuoGPT-o1
Chen et al. [20] addressed the challenge of enhancing reasoning in domain-specific applications by focusing on the medical field. In this domain, reliable reasoning is critical yet underexplored compared to mathematical tasks. The authors propose a novel two-stage training framework that leverages verifiable medical problems paired with a robust verifier to assess the correctness of model outputs. In the first stage, complex reasoning trajectories are generated and fine-tuned using a set of 40K carefully curated verifiable problems, ensuring each problem possesses a unique and verifiable solution. This stage is designed to guide the model through the intricacies of medical reasoning by using techniques such as backtracking, verification, and self-correction. The second stage involves reinforcement learning (using methods like PPO, outperforming alternatives such as RLOO and DPO) with verifier-based rewards, further refining the model’s reasoning capabilities. The resulting model, HuatuoGPT-o1, demonstrates substantial improvements over both general and medical-specific baselines, achieving an increase of approximately 80% on benchmarks like MedQA and PubMedQA. Additionally, experiments indicate that more complex reasoning chains (exceeding 700 tokens) yield better performance than straightforward reasoning approaches. Overall, the proposed model provides compelling evidence that a verifiable, two-stage training process can significantly enhance the complex reasoning abilities of LLMs in specialized domains such as medicine.
@@ III-G Search-o1
Li et al. [5] introduced Search-o1, an innovative framework designed to enhance the extended stepwise reasoning capabilities of large reasoning models (LRMs) by integrating a retrieval-augmented generation (RAG) mechanism with an advanced ”Reason-in-Documents” module. LRMs such as OpenAI-o1 exhibit impressive multi-step reasoning but are often hampered by knowledge gaps and uncertainties that can compromise the accuracy of their outputs. To mitigate these limitations, Search-o1 dynamically retrieves external knowledge during uncertain reasoning points using an agentic search workflow marked by specialized tokens to trigger search queries. The typically verbose and noisy retrieved documents are refined via the Reason-in-Documents module. This dedicated LLM call analyzes and distills the essential information before reintegrating it into the ongoing chain of thought. Extensive experiments across diverse domains, including science, mathematics, coding, and open-domain question answering, reveal that Search-o1 not only enhances overall accuracy—outperforming base models by 4.7% and traditional RAG systems by up to 29.6% on multi-hop questions—but also achieves higher performance than human experts in fields like physics and biology. This approach represents a significant step in creating more reliable and versatile intelligent systems by effectively filtering out noise and seamlessly merging retrieved knowledge with internal reasoning processes.
@@ III-H InternLM2
Cai et al. [44] The paper proposes InternLM2, a cutting-edge open-source large language model (LLM) that sets new standards in comprehensive evaluations across multiple dimensions and benchmarks. Developed by the Shanghai AI Laboratory, InternLM2 distinguishes itself through innovative pre-training and optimization techniques, effectively capturing long-term dependencies with an impressive context window of up to 128k tokens. The model’s pre-training process is meticulously detailed, encompassing diverse data types such as text, code, and specialized long-context data, and transitions from an initial 4k-token training phase to extended 32k-token stages, culminating in robust performance on challenging tests like the 200k ”Needle-in-a-Haystack” benchmark. Furthermore, InternLM2 is refined via Supervised Fine-Tuning (SFT) and a novel Conditional Online Reinforcement Learning from Human Feedback (COOL RLHF) strategy, which adeptly addresses issues like conflicting human preferences and reward hacking. Notably, the 8B parameter version of InternLM2 demonstrates remarkable efficiency, having been trained on only 4 trillion tokens—yielding over 75% in cost savings compared to similar models—while delivering superior results on complex reasoning tasks (scoring 83.0% on MATH-500 and 20.0% on AIME2024) and matching or surpassing contemporaries like Llama3.1-8B, Qwen2.5-7B, and OpenAI GPT-4o mini. This work marks a significant step forward in the open-source LLM landscape and provides valuable insights into model evolution and the practical realization of advanced reasoning capabilities in a cost-effective framework.
@@
Yao et al. [45] introduced MiniCPM-o 2.6, an advanced on-device multimodal large language model (MLLM) that extends the capabilities of its predecessors in the MiniCPM-V series. Uniquely designed to handle a wide array of modalities—including image, video, text, and audio—it delivers high-quality text and speech outputs in an end-to-end manner, all within an 8B parameter framework. Notably, MiniCPM-o 2.6 demonstrates leading visual capabilities, achieving an average score of 70.2 on OpenCompass and outperforming proprietary models like GPT-4o-202405, Gemini 1.5 Pro, and Claude 3.5 Sonnet in single and multi-image as well as video understanding tasks. Its state-of-the-art speech capabilities are evident in bilingual real-time conversations, enhanced ASR, speech-to-text translation performance, and flexible voice customization features such as emotion, speed, and style control. Furthermore, the model excels in multimodal live streaming, effectively processing continuous video and audio streams with real-time interaction, and exhibits robust OCR performance on images of varied aspect ratios and resolutions, surpassing models under 25B parameters on OCRBench. Another highlight is its superior efficiency—MiniCPM-o 2.6 achieves a remarkable token density, encoding 1.8 million pixels in only 640 tokens, thereby reducing latency and resource consumption.
@@ III-J KaLM-Embedding
Hu et al. [46] introduced KaLM-Embedding, a multilingual embedding model that emphasizes the importance of training data quality in the era of retrieval-augmented generation. Unlike prior work that often overlooks data curation, KaLM-Embedding is trained on a large corpus of cleaner, more diverse, and domain-specific data. The model leverages several key techniques to boost performance: persona-based synthetic data generation, which diversifies examples distilled from large language models; ranking consistency filtering, designed to eliminate noisy and uninformative samples; and semi-homogeneous task batch sampling to enhance training efficiency. Departing from traditional BERT-like architectures, the authors build on the Qwen2-0.5B pre-trained model, effectively adapting an auto-regressive language model framework to general embedding tasks. Extensive evaluations on the MTEB benchmark, spanning multiple languages, demonstrate that KaLM-Embedding outperforms comparable models with fewer than 1B parameters, achieving an average score of 64.53. Additionally, the model supports flexible embedding dimensions through its Matryoshka Representation Learning strategy, setting a new standard for multilingual embeddings in low-parameter regimes.
@@
Guan et al. [19] introduced rStar-Math, a framework that leverages small language models (SLMs) to achieve state-of-the-art performance in mathematical reasoning, rivaling and even surpassing the capabilities of larger models like OpenAI o1-preview. The approach employs deep thinking via Monte Carlo Tree Search (MCTS), where an SLM-based policy model generates multiple code-augmented chain-of-thought (CoT) reasoning trajectories. Each step in these trajectories comprises a natural language explanation and executable Python code, ensuring that reasoning steps are verified through code execution. A key innovation lies in the novel synthesis of CoT data via extensive MCTS rollouts, which produce step-by-step verified reasoning paths to train the policy model. Complementing this, the authors propose an improved process reward model (PRM) training method that avoids naive step-level scoring by leveraging binary code verification outcomes, and they introduce a self-evolution strategy whereby both the policy SLM and PRM are iteratively refined from an initial dataset of 747k math problems. After four rounds of self-evolution involving millions of synthesized solutions, rStar-Math boosts accuracy on the MATH benchmark to 90.0% using a 7B parameter model, outperforming o1-preview, and solves 53.3% of problems on the USA Math Olympiad (AIME), ranking within the top 20% among high school competitors.
@@
Microsoft [47] introduced phi-4, a 14-billion parameter language model that significantly advances performance on STEM-focused QA tasks by leveraging a training strategy centered on data quality. Unlike many models that rely predominantly on organic data sources such as web content or code, phi-4 integrates large-scale, high-quality synthetic data throughout its pre-training process. This synthetic data is generated through multi-agent, self-revision workflows, enabling phi-4 to distill and surpass the capabilities of its teacher model, GPT-4, particularly in reasoning, mathematics, and code generation. Despite minimal modifications to the phi-3 architecture, phi-4, trained on 9.8 trillion tokens and supporting a 16K token context length, performs comparably to models with 70B parameters. The model was developed with a comprehensive post-training scheme—including supervised fine-tuning (SFT) and Direct Preference Optimization (DPO) for safety alignment—and required 21 days of training on 1920 H100-80G GPUs. Phi-4 exemplifies how strategic data generation and refined training curricula can dramatically enhance model performance, especially in specialized STEM domains.
@@ III-M DeepSeek-V3
DeepSeek Team proposed [48] DeepSeek-V3, a state-of-the-art Mixture-of-Experts (MoE) language model comprising 671 billion parameters, with 37 billion activated per token during inference. DeepSeek-V3 leverages advanced architectures such as Multi-head Latent Attention (MLA) and DeepSeekMoE to ensure efficient inference and cost-effective training, previously validated in its predecessor, DeepSeek-V2. Notably, the model pioneers an auxiliary-loss-free strategy for load balancing and adopts a multi-token prediction training objective to boost performance. Pre-training was conducted on an expansive corpus of 14.8 trillion high-quality tokens, followed by rigorous Supervised Fine-Tuning and Reinforcement Learning stages, enabling DeepSeek-V3 to achieve performance that matches or exceeds leading closed-source models like OpenAI’s GPT-4o and Anthropic’s Claude-Sonnet-3.5. Despite its scale, the training process was remarkably stable and efficient, requiring only 2.788 million H800 GPU hours—approximately \$5.5 million in cost—while supporting FP8 training and achieving a throughput of 60 tokens per second. Additionally, the model offers significant cost advantages with low per-input and per-output expenses, and it is trained on both English and Chinese data, marking a notable engineering milestone in large-scale language models.

### v2

@@ III-C Mistral AI Small 3 24B
Mistral Small 3 [mistral_small_3] is a highly efficient language model, competing with Llama 3.3 70B and Qwen 32B, while being 3× faster on the same hardware. Optimized for 80% of generative AI tasks, it excels in instruction-following with low latency and achieves 81% accuracy on MMLU at 150 tokens/s. Fewer layers significantly reduce inference time, making it ideal for local deployment. Unlike models such as DeepSeek R1, Mistral Small 3 has not been trained using Reinforcement Learning (RL) or synthetic data, making it an earlier-stage model in the production pipeline with the potential for additional refinement.
Figure 3: Training Pipelines and Model Variants of DeepSeek-R1 and R1-Zero.
@@ III-D DeepSeek-R1
The DeepSeek-R1 [deepseekai2025] series proposes advanced reasoning-optimized language models, including DeepSeek-R1-Zero, which is trained exclusively via Reinforcement Learning (RL) without Supervised Fine-Tuning (SFT), and DeepSeek-R1, which incorporates multi-stage training with cold-start data for improved reasoning and language coherence. DeepSeek-R1 matches OpenAI’s o1-1217 in reasoning tasks, leveraging key architectural innovations: Mixture of Experts (MoE) for efficient parameter utilization, Multihead Latent Attention (MLA) for reduced memory and compute costs, Multi-Token Prediction (MTP) for 2-3x faster token generation, and FP8 Quantization, which cuts memory usage by 75% while maintaining stability. DeepSeek has released DeepSeek-R1-Zero, DeepSeek-R1, and six dense models (1.5B–70B) to support open research, providing a scalable, high-performance foundation for AI-driven reasoning and inference.
In January 2025, DeepSeek introduced two influential models, DeepSeek-R1-Zero and DeepSeek-R1, which stood out for their impressive reasoning capabilities achieved at minimal training costs. These models challenged the notion that advanced reasoning requires extensive computational resources, highlighting a new direction in efficient AI development. As presented in Fig. 3 (a), DeepSeek-R1-Zero is built directly from the DeepSeek-V3-Base model without any supervised fine-tuning. Instead, it uses a rule-based reinforcement learning method called Group Relative Policy Optimization (GRPO), which uses structured reasoning rules to guide training. This design enables a simpler and more scalable pipeline. DeepSeek-R1 builds upon this foundation by addressing R1-Zero’s limitations, such as low output fluency and language inconsistency. It incorporates carefully selected high-quality data and multiple training stages, including reinforcement learning with rejection sampling, to enhance its general-purpose abilities and improve alignment with human preferences as presented in Fig. 3 (b).
To ensure accessibility across various platforms, DeepSeek also released a family of distilled models based on DeepSeek-R1, ranging from 1.5 billion to 70 billion parameters as presented in Fig. 3 (c). These smaller models are fine-tuned on synthetic data generated by the original R1 model, making them lightweight yet capable of strong reasoning performance. This approach allows the models to run efficiently on devices with limited hardware capabilities. Across multiple benchmarks in math, coding, general knowledge, and writing, DeepSeek-R1 models show strong competitiveness while offering significant cost advantages over other commercial options. In some cases, the models are up to 20 to 50 times more cost-effective than counterparts like OpenAI’s o1, making them a practical solution for research and real-world applications.
@@
Qwen2.5-VL [qwen2.5_vl] is a significant upgrade to Qwen2-VL [wang2024], enhancing vision-language capabilities across three sizes (3B, 7B, 72B). The Qwen2-VL Series proposes Naive Dynamic Resolution (NDR) for adaptive visual tokenization, enhancing efficiency and accuracy in vision-language tasks. It integrates Multimodal Rotary Position Embedding (M-RoPE) for improved text-image-video fusion and follows a unified image-video processing paradigm. By scaling both model size (2B, 8B, 72B) and training data, Qwen2-VL achieves competitive performance, with Qwen2-VL-72B matching GPT-4o and Claude 3.5-Sonnet on multimodal benchmarks. These innovations position Qwen2-VL as a state-of-the-art LVLM, advancing vision-language integration.
@@ III-F HuatuoGPT-o1
Chen et al. [chen2024huatuogpt] addressed the challenge of enhancing reasoning in domain-specific applications by focusing on the medical field. In this domain, reliable reasoning is critical yet underexplored compared to mathematical tasks. The authors propose a novel two-stage training framework that leverages verifiable medical problems paired with a robust verifier to assess the correctness of model outputs. In the first stage, complex reasoning trajectories are generated and fine-tuned using a set of 40K carefully curated verifiable problems, ensuring each problem possesses a unique and verifiable solution. This stage is designed to guide the model through the intricacies of medical reasoning by using techniques such as backtracking, verification, and self-correction. The second stage involves reinforcement learning (using methods like PPO, outperforming alternatives such as RLOO and DPO) with verifier-based rewards, further refining the model’s reasoning capabilities. The resulting model, HuatuoGPT-o1, demonstrates substantial improvements over both general and medical-specific baselines, achieving an increase of approximately 80% on benchmarks like MedQA and PubMedQA. Additionally, experiments indicate that more complex reasoning chains (exceeding 700 tokens) yield better performance than straightforward reasoning approaches. Overall, the proposed model provides compelling evidence that a verifiable, two-stage training process can significantly enhance the complex reasoning abilities of LLMs in specialized domains such as medicine.
@@ III-G Search-o1
Li et al. [li2025search] introduced Search-o1, an innovative framework designed to enhance the extended stepwise reasoning capabilities of large reasoning models (LRMs) by integrating a retrieval-augmented generation (RAG) mechanism with an advanced ”Reason-in-Documents” module. LRMs such as OpenAI-o1 exhibit impressive multi-step reasoning but are often hampered by knowledge gaps and uncertainties that can compromise the accuracy of their outputs. To mitigate these limitations, Search-o1 dynamically retrieves external knowledge during uncertain reasoning points using an agentic search workflow marked by specialized tokens to trigger search queries. The typically verbose and noisy retrieved documents are refined via the Reason-in-Documents module. This dedicated LLM call analyzes and distills the essential information before reintegrating it into the ongoing chain of thought. Extensive experiments across diverse domains, including science, mathematics, coding, and open-domain question answering, reveal that Search-o1 not only enhances overall accuracy outperforming base models by 4.7% and traditional RAG systems by up to 29.6% on multi-hop questions but also achieves higher performance than human experts in fields like physics and biology. This approach represents a significant step in creating more reliable and versatile intelligent systems by effectively filtering out noise and seamlessly merging retrieved knowledge with internal reasoning processes.
@@ III-H InternLM2
Cai et al. [cai2024internlm2] propose InternLM2, a cutting-edge open-source large language model (LLM) that sets new standards in comprehensive evaluations across multiple dimensions and benchmarks. Developed by the Shanghai AI Laboratory, InternLM2 distinguishes itself through innovative pre-training and optimization techniques, effectively capturing long-term dependencies with an impressive context window of up to 128k tokens. The model’s pre-training process is meticulously detailed, encompassing diverse data types such as text, code, and specialized long-context data, and transitions from an initial 4k-token training phase to extended 32k-token stages, culminating in robust performance on challenging tests like the 200k ”Needle-in-a-Haystack” benchmark. Furthermore, InternLM2 is refined via Supervised Fine-Tuning (SFT) and a novel Conditional Online Reinforcement Learning from Human Feedback (COOL RLHF) strategy, which adeptly addresses issues like conflicting human preferences and reward hacking. Notably, the 8B parameter version of InternLM2 demonstrates remarkable efficiency, having been trained on only 4 trillion tokens yielding over 75% in cost savings compared to similar models while delivering superior results on complex reasoning tasks (scoring 83.0% on MATH-500 and 20.0% on AIME2024) and matching or surpassing contemporaries like Llama3.1-8B, Qwen2.5-7B, and OpenAI GPT-4o mini. This work marks a significant step forward in the open-source LLM landscape and provides valuable insights into model evolution and the practical realization of advanced reasoning capabilities in a cost-effective framework.
@@
Yao et al. [yao2024minicpm] introduced MiniCPM-o 2.6, an advanced on-device multimodal large language model (MLLM) that extends the capabilities of its predecessors in the MiniCPM-V series. Uniquely designed to handle a wide array of modalities, including image, video, text, and audio, it delivers high-quality text and speech outputs in an end-to-end manner, all within an 8-bit parameter framework. Notably, MiniCPM-o 2.6 demonstrates leading visual capabilities, achieving an average score of 70.2 on OpenCompass and outperforming proprietary models like GPT-4o-202405, Gemini 1.5 Pro, and Claude 3.5 Sonnet in single and multi-image as well as video understanding tasks. Its state-of-the-art speech capabilities are evident in bilingual real-time conversations, enhanced ASR, speech-to-text translation performance, and flexible voice customization features such as emotion, speed, and style control. Furthermore, the model excels in multimodal live streaming, effectively processing continuous video and audio streams with real-time interaction, and exhibits robust OCR performance on images of varied aspect ratios and resolutions, surpassing models under 25B parameters on OCRBench. Another highlight is its superior efficiency MiniCPM-o 2.6 achieves a remarkable token density, encoding 1.8 million pixels in only 640 tokens, thereby reducing latency and resource consumption.
@@ III-J KaLM-Embedding
Hu et al. [hu2025kalm] introduced KaLM-Embedding, a multilingual embedding model that emphasizes the importance of training data quality in the era of retrieval-augmented generation. Unlike prior work that often overlooks data curation, KaLM-Embedding is trained on a large corpus of cleaner, more diverse, and domain-specific data. The model leverages several key techniques to boost performance: persona-based synthetic data generation, which diversifies examples distilled from large language models; ranking consistency filtering, designed to eliminate noisy and uninformative samples; and semi-homogeneous task batch sampling to enhance training efficiency. Departing from traditional BERT-like architectures, the authors build on the Qwen2-0.5B pre-trained model, effectively adapting an auto-regressive language model framework to general embedding tasks. Extensive evaluations on the MTEB benchmark, spanning multiple languages, demonstrate that KaLM-Embedding outperforms comparable models with fewer than 1B parameters, achieving an average score of 64.53. Additionally, the model supports flexible embedding dimensions through its Matryoshka Representation Learning strategy, setting a new standard for multilingual embeddings in low-parameter regimes.
@@
Guan et al. [guan2025rstar] introduced rStar-Math, a framework that leverages small language models (SLMs) to achieve state-of-the-art performance in mathematical reasoning, rivaling and even surpassing the capabilities of larger models like OpenAI o1-preview. The approach employs deep thinking via Monte Carlo Tree Search (MCTS), where an SLM-based policy model generates multiple code-augmented chain-of-thought (CoT) reasoning trajectories. Each step in these trajectories comprises a natural language explanation and executable Python code, ensuring that reasoning steps are verified through code execution. A key innovation lies in the novel synthesis of CoT data via extensive MCTS rollouts, which produce step-by-step verified reasoning paths to train the policy model. Complementing this, the authors propose an improved process reward model (PRM) training method that avoids naive step-level scoring by leveraging binary code verification outcomes, and they introduce a self-evolution strategy whereby both the policy SLM and PRM are iteratively refined from an initial dataset of 747k math problems. After four rounds of self-evolution involving millions of synthesized solutions, rStar-Math boosts accuracy on the MATH benchmark to 90.0% using a 7B parameter model, outperforming o1-preview, and solves 53.3% of problems on the USA Math Olympiad (AIME), ranking within the top 20% among high school competitors.
@@
Microsoft [abdin2024phi] introduced phi-4, a 14-billion parameter language model that significantly advances performance on STEM-focused QA tasks by leveraging a training strategy centered on data quality. Unlike many models that rely predominantly on organic data sources such as web content or code, phi-4 integrates large-scale, high-quality synthetic data throughout its pre-training process. This synthetic data is generated through multi-agent, self-revision workflows, enabling Phi-4 to distill and surpass the capabilities of its teacher model, GPT-4, particularly in reasoning, mathematics, and code generation. Despite minimal modifications to the phi-3 architecture, phi-4, trained on 9.8 trillion tokens and supporting a 16K token context length, performs comparably to models with 70B parameters. The model was developed with a comprehensive post-training scheme including supervised fine-tuning (SFT) and Direct Preference Optimization (DPO) for safety alignment and required 21 days of training on 1920 H100-80G GPUs. Phi-4 exemplifies how strategic data generation and refined training curricula can dramatically enhance model performance, especially in specialized STEM domains.
@@ III-M DeepSeek-V3
DeepSeek Team proposed [liu2024deepseek] DeepSeek-V3, a state-of-the-art Mixture-of-Experts (MoE) language model comprising 671 billion parameters, with 37 billion activated per token during inference. DeepSeek-V3 leverages advanced architectures such as Multi-head Latent Attention (MLA) and DeepSeekMoE to ensure efficient inference and cost-effective training, previously validated in its predecessor, DeepSeek-V2. Notably, the model pioneers an auxiliary-loss-free strategy for load balancing and adopts a multi-token prediction training objective to boost performance. Pre-training was conducted on an expansive corpus of 14.8 trillion high-quality tokens, followed by rigorous Supervised Fine-Tuning and Reinforcement Learning stages, enabling DeepSeek-V3 to achieve performance that matches or exceeds leading closed-source models like OpenAI’s GPT-4o and Anthropic’s Claude-Sonnet-3.5. Despite its scale, the training process was remarkably stable and efficient, requiring only 2.788 million H800 GPU hours, approximately \$5.5 million in cost, while supporting FP8 training and achieving a throughput of 60 tokens per second. Additionally, the model offers significant cost advantages with low per-input and per-output expenses, and it is trained on both English and Chinese data, marking a notable engineering milestone in large-scale language models.